Consider the following LPP: Maximize z = 50x1 + 20x2 + 30x3 subject to 2x1 + x2 + 3x3 + 90 (Resource A) x1 + 2x2 + x3 + 50 (Resource B) x1 + x2 + x3 + 80 (Resource C) x1, x2 , x3 > 0 The final simplex table is Basis cj  x1 x2 x3 s1 s2 s3 Solution 50 20 30 0 0 0 x1 50 1 -1 0 1 -1 0 40 x3 30 0 3 1 -1 2 0 10 s3 0 0 -1 0 0 -1 1 30 zj 50 40 30 20 10 0 2300 cj-zj 0 -20 0 -20 -10 0 a. What is the range of optimality for the contribution rate of the variable x1? b. Obtain the range of feasibility for the Resource A.
Answer #1

a) Range of optimality for the contribution rate of variable x1 is following:

Lower limit = 50-10 = 40

Upper limit = 50+infinity = infinity

b) Range of feasibility for resource A is following:

Lower limit = 90-90 = 0

Upper limit = 90+10 = 100